Use the following table to answer questions 10 – 13:

# Units Produced

Total Revenue

Total Costs

0

0

0

1

100

50

2

180

110

3

250

180

4

290

270

5

310

380

10. What is the marginal revenue of producing the 2nd unit?

a) 50.

b) 70.

c) 90.

d) 80.

11. What is the marginal cost of producing the 4th unit?

a) 270.

b) 110.

c) 50.

d) 90.

12. At what level of output does marginal cost equal marginal revenue?

a) 1.

b) 2.

c) 3.

d) 4.

13. What is the level of net benefits when three units are produced?

a) 120.

b) 70.

c) -70.

d) 20.
Answer #1

10. d) 80

Marginal revenue = 180-100= 80

11. d) 90

Marginal cost = 270 - 180 = 90

12. c) 3

Marginal revenue at Q= 3 is 250-180= 70

Marginal cost at Q= 3 is 180-110= 70

13. b) 70
Net benefits at Q= 3 is 250-180= 70